I have to solve the following system by addition:
5x - 3y = 13
4x - 3y = 11
I have tried using the example from our text and I come out with x=2, y= -1 but when Y check these #'s into the equations they do not work. Please help.

Answer by funmath(2933)

You're right you must be checking it wrong:
5(2)-3(-1)=13
10+3=13
13=13
The first equation checks, now check the second equation:
4(2)-3(-1)=11
8+3=11
11=11
The second equation checks.
Congratulations, you were right!!!!